OpenStudy (anonymous):

how do you graph y+2=1/3 (x+2)

OpenStudy (kc_kennylau):

https://www.desmos.com/calculator/2ttmlj0jhe

OpenStudy (anonymous):

Change it to slope-intercept form. y = mx + b B - y intercept m - slope Here is the first step: \[\huge 3 * (y + 2) = \frac{ (x+2) }{ \cancel {3} } * \cancel{3}\]

OpenStudy (anonymous):

Simplify it and convert it to slope-intercept form, once you do that we'll go from there..

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hi, The first step is to put it in Slope-Intercept Form! y = mx + b. Where b is my y-intercept and m is my slope. y+2=1/3 (x+2) y = 1/3x (I got this by subtracting 2 from both sides, because I want to solve for y) Now, I want to use rise-over-run on my graph. So, here, we have 1/3. rise/run. I rise 1 space, and go over 3 spaces! I do this and put a dot. From that dot I do it even more! When I got two dots, I can draw a line! :)  I hope this helped! Please - come back to OpenStudy, and if you EVER have a problem, feel free to message or tag a moderator (purple name guys!) or Ambassador (Purple A guys, like me!) OpenStudy Ambassador; Compassionate.

OpenStudy (anonymous):

goformittens for 2k14